What is the maximum area of a triangle having 15cm as the sum of its base and height ? Please Help :)
Answer by scott8148(6628)

b+h=15 ___ b=15-h

area=bh/2 ___ a=(15-h)(h)/2 ___ a=(15/2)h-(1/2)h^2

maximum area is on the axis of symmetry ___ h=(-15/2)/(2(-1/2)) ___ h=15/2

b=15-h ___ b=15-(15/2) ___ b=15/2

a=(15/2)(15/2)/2 ___ a=225/8